OpenStudy (anonymous):

Its a complex question.. Whe have an ellipse that crosses trough the next 5 points: A(0,4) B(6,2) C(2,-2) D(-2,-2) E(-6,2) Find the ellipse formula

OpenStudy (anonymous):

It has many solutions. I will let you know why.

OpenStudy (anonymous):

The general equations of a conic is of the form \[ g(x,y)=a x^2+b x y+c y^2+d x+e y+f=0 \] You have to find six unknowns and you have five information, you can take one unknown arbitrary. I took f=112 to get nice answer.

OpenStudy (anonymous):

The five equations are obtained by putting g (each point) =0 You will get five equations \[ 16 c+4 e+112=0\\ 36 a+12 b+4 c+6 d+2 e+112=0\\ 4 a-4 b+4 c+2 d-2e+112=0\\ 4 a+4 b+4 c-2 d-2 e+112=0 \\ 36 a-12 b+4 c-6 d+2 e+112=0 \] You put them in wolframalpha and you solve

OpenStudy (anonymous):

Actually, I used Mathematica to solve You find {a= -3, b= 0, c = -13, d= 0, e= 24}

OpenStudy (anonymous):

Your equation will be \[ -3 x^2-13 y^2+24 y+112=0 \]

OpenStudy (anonymous):

Here is the ellipse with your points on it @ganeshie8 @Loser66 @ikram002p

ganeshie8 (ganeshie8):

yess.. just wondering.. if we limit to ellipses whose major/minor axis are parallel to xy axes, then we may get an unique solution...

ganeshie8 (ganeshie8):

like the one in ur graph :)

ganeshie8 (ganeshie8):

oh its the oly ellipse possible if we limit to translation... all other ellipses wud require rotation !

OpenStudy (anonymous):

Actually, @ganeshie8 you might be right. Try it

OpenStudy (anonymous):

Actually, I tried f=2 and I go the same ellipse.

OpenStudy (ikram002p):

it said in the qs its ellipse so i guss its oly onr sol , we just need to match the hypo. wid ellipse fomulla

OpenStudy (ikram002p):

we cud use the points B(6,2) ,E(-6,2) C(2,-2), D(-2,-2) such that f'(6)=-f'(-6) f'(2)=-f'(-2) hummm but im not sure it wont work :|

OpenStudy (ikram002p):

i think ya its complex but it can be solved

OpenStudy (ikram002p):

if u cud solve the equation 1,2,3,4,5

OpenStudy (anonymous):

actually, i was searching for the method. The easy way out would have been to just use mathematica or geogebra.

OpenStudy (ikram002p):

just see eliasab sol !